OpenStudy (chillhill):

Find x if G(x) = -3 G(x)= (2x+3)/(x-4)

OpenStudy (jango_in_dtown):

(2x+3)/(x-4)=-3 implies 2x+3=-3x+12 implies 5x=9 implies x=9/5